(e) The probability that a programmer will not complete writing the code is 0.12. Given that there were 25 programmers, find the probability that not more than 5 of these programmers will complete writing the code..
Answer #1

X : number of programmers who will not complete writing the code.

X ~ bin (n,p)

where, n = 25

p = 0.12

q = (1-p) = (1-0.12)= 0.88

the pmf of the distribution be:-

the probability that not more than 5 of these programmers will complete writing the code be:-

***in case of doubt, comment below. And if u liked the solution, please like.